OpenStudy (anonymous):

Which statement is true? f(x)=2√(x) and g(x)=x^2-1

OpenStudy (skullpatrol):

What do you mean by true?

hartnn (hartnn):

options are missing...

OpenStudy (hba):

@Katieann16 Please respond....

OpenStudy (anonymous):

A: (fog)(x)=2(x^2-1) B: (fog)(x)=2x(x^2-1) C: (gof)(x)=4x^2-1 D: (gof)(x)=4x-1

OpenStudy (anonymous):

sorry guys

OpenStudy (hba):

Thanks for replying.

OpenStudy (hba):

Do you know how to find (fog)x ?

hartnn (hartnn):

to find (fog)(x), you replace every 'x' in f(x) by g(x).

OpenStudy (anonymous):

yes

OpenStudy (hba):

Let me show you an example f(x)=x^2-1 and g(x)=x+2 (fog)x=(x+2)^2-1

OpenStudy (hba):

@Katieann16 Why don't you try solving it ?

OpenStudy (hba):

@hartnn I am sorry i gotta leave,you can proceed.

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

to find (fog)(x), you replace every 'x' in f(x) by g(x). so, find (fog)(x) and tell me what u get.

OpenStudy (anonymous):

so every f(x) is g(x)?

hartnn (hartnn):

? no.... ok, try finding, (gof) (x) you have the function g(x), wherever u see 'x' in it replace it by 2√(x) and tell me what u get for (gof)(x).

OpenStudy (anonymous):

im confused .

hartnn (hartnn):